Nutrition Calculations

1. One cheeseburger has 340 kcal and 25 grams of fat. What percentage of total kcal comes from fat? (round to the nearest whole number)

2. A peanut-butter granola bar contains 2 grams of protein, 8 grams of fat, and 20 grams of carbohydrate. How many kcal in this bar come from fat?

3. Lee wants to gain 10lbs of lean muscle mass. His dietitian suggested he consume 2400 kcal each day and that 25% of his kcal come from protein. How many kcal from protein should he consume each day? (round to the nearest whole number)

4. A 3-ounce chicken breast contains 21 grams of protein. How many kcals in the 3-ounce chicken breast come from protein?

5. Jan was recently diagnosed with diabetes. Her dietitian suggested she should consume 1500 calories a day and 55% of her calories should come from carbohydrate. How many grams of carbohydrate should she consume each day? (round to the nearest whole number)

please show me how to get the answer and the answer
Answer #1

1. One cheeseburger has 340 Kcal.

one cheeseburger has 25 grams of fat. one gram of fat has 9 cal.

So, 25 gm fat has (25*9) cal= 225 cal= 225/1000 Kcal. = 0.225 Kcal (As, 1 Kcal= 1000 cal.)

Thus, percentage of total Kcal from fat= (0.225/340)*100= 0.06618%

2. A peenut butter grenola bar contains 2 gm of protein, 8 gm of fat and 20 gm of carbohydrate.

One gram fat contains 9 cal and 1 Kcal= 1000 cal.

So, the total Kcal in this bar that come from fat= 8*9/1000 Kcal = 0.072 Kcal

3. Lee needs to consume 2400 Kcal per day. And, 25% of his Kcal must come from protein.

So, the total Kcal of protein that he must consume= 2400*(25/100) Kcal= 600 Kcal.

4. A 3 ounce chicken breast contains 21 gms of protein.

One gram of protein contains 4 cal.

So, the total Kcal obtained from proteins in chicken breast= (4*21)/1000 Kcal= 0.0084 Kcal.

5. Jan must consume 1500 calories and 55% of her calories must come from carbohydrate.

So, Total Kcal Jan must consume from carbohydrate= 1500*(55/100) cal = 825 cal

One gram carbohydrate has 4 calories.

So, the amount of carbohydrate Jan needs to consume= 825/4 gm= 206.25 gm

Thank You.

Happy Chegging!